You are given the data below for 2008 for the imaginary country of Amagre, whose currency is the G.

Consumption 350 billion G
Transfer payments 100 billion G
Investment 100 billion G
Government purchases 200 billion G
Exports 50 billion G
Imports 150 billion G
Bond purchases 200 billion G
Earnings on foreign investments 75 billion G
Foreign earnings on Amagre investment 25 billion G

Compute net foreign investment.
Compute net exports.
Compute GDP.
Compute GNP.

In addition to responding with a quantitative answer, briefly, describe how you arrived at your answers.
1
Expert's answer
2016-07-27T08:13:02-0400
Net foreign investment is equal to net exports. The value of exports minus imports is net exports.
Net exports = Exports – Imports = 50 – 150 = -100 billion G.
As a result net foreign investment equals 100 billion G.
Gross domestic product is the value of finished goods and services manufactured within a country in a defined period of time. GDP equals to the sum of consumption, government expenditures, investments and net exports.
GDP = C + I + G + NX
GDP = 350 + 100 + 200 – 100 = 550 billion G.
Gross national product equals GDP plus income earnings from foreigners minus income payments to other countries.
GNP = C + I + G + NX + NY
GNP = GDP + NY
GNP = 550 + (25 – 75) = 500 billion G.